OpenStudy (anonymous):

Solve the problem. Round to the nearest hundredth of a percent if needed. During clinical trials of a new drug intended to reduce the risk of heart attack, the following data indicate the occurrence of adverse reactions among 1100 adult male trial members. What is the probability that an adult male using the drug will experience nausea? I will post table

OpenStudy (anonymous):

A. 2.02% B. 1.73% C. 27.59% D. 2.18%

jimthompson5910 (jim_thompson5910):

How many adults experience nausea?

jimthompson5910 (jim_thompson5910):

According to the table

OpenStudy (anonymous):

24

jimthompson5910 (jim_thompson5910):

So probability of getting nausea (as a percentage) is (24/1100)*100

jimthompson5910 (jim_thompson5910):

since there are 1100 people total

jimthompson5910 (jim_thompson5910):

the portion "*100" converts the decimal to a percent

OpenStudy (anonymous):

i thought you have to add all the numbers in the table then divide 24/x x being the numbers added..

jimthompson5910 (jim_thompson5910):

If you add up all the numbers, you get 87 But there are 1100 people total...so something is off

jimthompson5910 (jim_thompson5910):

You should add up all the numbers to get the reported total of 1100

OpenStudy (anonymous):

ok so it is D?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

I'm getting D as well (assuming it's 1100 people total...and it's not a typo)

OpenStudy (anonymous):

ok thanks

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

could you just help me in 2 more plz?

jimthompson5910 (jim_thompson5910):

sure thing

OpenStudy (anonymous):

A game spinner has regions that are numbered 1 through 9. If the spinner is used twice, what is the probability that the first number is a 3 and the second is a 6? A. 1/18 B. 1/81 C. 1/9 D. 2/3 Im getting B i just want to make sure it is right

jimthompson5910 (jim_thompson5910):

correct, (1/9)*(1/9) = 1/81 is your answer

OpenStudy (anonymous):

thanks! This is the last one! Solve the problem. Round to the nearest dollar if needed. Looking ahead to retirement, you sign up for automatic savings in a fixed-income 401K plan that pays 5% per year compounded annually. You plan to invest $3500 at the end of each year for the next 15 years. How much will your account have in it at the end of 15 years? A. $77,295 B. $75,525 C. $76,823 D. $73,982

jimthompson5910 (jim_thompson5910):

this is an annuity problem, one sec while I find the formula

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

FV = P*((1+r)^n - 1)/r FV = 3500*((1+0.05)^15 - 1)/0.05 FV = 75,524.9725587959 FV = 75,525 (round to the nearest dollar)

jimthompson5910 (jim_thompson5910):

FV stands for future value

OpenStudy (anonymous):

thank you so much! I really appreciate it!

jimthompson5910 (jim_thompson5910):

sure thing